# Mlynská dolina, 842 48, Bratislava, Slovakia

### v1

@@ 3.3 Numerical results
In this section, we gather the results of numerical simulations performed on a personal computer. The reference energies denoted $\infty$ as $M\rightarrow\infty$ corresponds to the limit of ordinary space are taken from aquino2007highly . The results provided are eigenvalues of the matrix defined in (11), showing different numbers of layers. The energies are in hartrees and the distances in bohrs. The large matrix values shown in table 3 were obtained in 10 hours on a personal computer, and the code was written in C++. We can observe that above the matrix size of the order of $10^{4}$, numerical errors seem to be piling up.
Figure 2: Radial parts of the probability distribution for all energy eigenstates for the model with $r_{0}=10$ and $M=100$. The same plot is obtained — in reverse order, starting from the largest eigenvalue, see (14) — for the same model but with repulsive potential. Green lines show numerical solutions of (11), and red lines show exact solutions for an unconfined atom.
@@
Figure 3: Comparison of the radial part of our solution using the matrix defined (11) and the ordinary-space solution for confided hydrogen atom from reference aquino2007highly for $r_{0}=10$ and 2S orbital. We have also included the unconfined solution to show the effect of the barrier.
Figure 4: Comparison of the radial part of our solution using the matrix defined (11) and the ordinary-space solution for confided hydrogen atom from reference aquino2007highly for $r_{0}=20$ and $4$S orbital. We have also included the unconfined solution to show the effect of the barrier.
$M$ $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ reference
Table 3: Values of the ground state energy for $r_{0}=10$ and different number of layers. The largest value took approximately $10$ hours of computational time on a personal computer. The reference is taken from aquino2007highly .
$r_{0}\backslash M$infty as a reference 100 1000 10000
Table 4: The 1S orbital,
@@
Table 5: $E^{II}$ are the largest eigenvalues for repulsive potential, $E^{I}$ are the smallest eigenvalues for the attractive potential. The next column shows the theoretical prediction, and the last one compares this with obtained numerical values. Energies are in hartree.
Figure 5: Comparison of the eigenvectors for the third smallest eigenvalue in the case of attractive potential and the third largest eigenvalue for the repulsive potential, these states have dual energy according to (14). These values were obtained numerically for $M=1000$, $r_{0}=40$ and $n=3$.
$r_{0}\backslash M$infty as a reference 100 1000 10000
20 -0.124987114312918 -0.124677720985899 -0.$1$0.124987102906836
Table 6: The 2S orbital.
$r_{0}\backslash M$infty as a reference 100 1000 10000
Table 7: The 3S orbital.
$r_{0}\backslash M$infty as a reference 100 1000 10000
Table 8: The 2P orbital.
$r_{0}\backslash M$infty as a reference 100 1000 10000
Table 9: The 3P orbital.
@@ 4 Conclusions and outlook
We have demonstrated a simple model of quantum space that can be used in a certain limit for a description of ordinary physics, for example, the hydrogen atom problem. The defining feature of this model is that it is expressed in terms of matrices and can, therefore, easily be evaluated on a computer.
To showcase the behaviour of the model, we have analysed the behaviour of hydrogen confined in a spherical cavity and compared our results to those previously obtained in high-precision studies; the agreement is up to numeral decimal places. We have set up a Hamiltonian expressed in terms of a matrix; its eigenvalues form the energy spectrum and the corresponding eigenvectors for the radial dependency. The model can be easily amended to have a different form of potential, making this a rather flexible model.
For the problem we analyzed, it was possible to take only a single value $(l,m)$; however, angular-dependant problems can be analyzed in the defined setting as well; one just needs to use the full form of the matrix $\Psi$ and the Hamiltonian.
In our case, we sent the constant describing the quantumness of space, $\lambda$, to 00 to simulate the physics in ordinary space. However, one can be motivated to choose a nonzero value to describe a space with a granular structure or to capture the essence of a finite-length scale feature of the studied model.
We have verified the behaviour of the confined particle with respect to the different radii of the cavity and showed how the particle is affected in certain regimes. We have also studied the behaviour in the strong quantum-space regime. The first important observation was that increasing the value of $\lambda$ pushes the particle away from the centre; this means that the quantumness of space manifests as an outward pressure. This is important, for example, to understand the behaviour of regularized singularity in studies of microscopic black holes sensitive to this length scale Nicolini:2005vd ; Kovacik:2021qms . In addition to that pgk15 , it has been shown that a new form of bounded states should appear for ultra-high energies. We have shown that these states indeed appear and confirm the pattern of mirroring energies. Since our parameter, $\lambda$, can also be understood as describing a length scale that is not of quantum-space origin, but perhaps due to two granular features of the material, it might be interesting to consider whether such spurious states might be experimentally observable.
Also, our plan is to investigate the behaviour of a free particle in moving confinement, studying the pressure it produces (in the strong quantum space limit); those studies should provide some insight into high-density regimes of the theory of quantum space, which are needed for the description of the earliest stages or the universe or the behaviour of black hole singularities as described in pgk15 ; pg13 .
@@ Acknowledgements.
This research was supported by VEGA 1/0025/23 grant Matrix models and quantum gravity and MUNI Award for Science and Humanities funded by the Grant Agency of Masaryk University. The authors also thank the hospitality of the Physikzentrum Bad Honnef, where the school co-funded and co-organised by COST (European Cooperation in Science and Technology) through the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG); the work in this paper was initiated there. We would like to thank Lukáš Konečný for his comments.
Appendix A Explicit construction of the $su(2)$ representation with Hermitian matrices of rank $N$
Let us begin with the construction of the angular momentum generators satisfying $[L^{i},L^{j}]=i\varepsilon_{ijk}L^{k}$. We can choose a basis for which eigenstates of $L_{3}$ are unit vectors of the form $e_{l-m}=(0,\ldots,0,1,0,\ldots,0)$ such that $L_{3}e_{m}=me_{m}$. Then, we define two matrices that increase this value by $1$, that is $L_{\pm}e_{m}\sim e_{m\pm 1}$. Taking into account the normalisation, one obtains
@@ Appendix A Explicit construction of the $su(2)$ representation with Hermitian matrices of rank $N$
and $L_{-}=L_{+}^{T}$. The diagonal coefficients are $\sqrt{\left(l+1-m\right)m}$. Now one can take
@@ Appendix B Radial term on the fuzzy onion
${\cal{K}}=\lambda^{-2}\begin{pmatrix}-2&\frac{2}{1}&0&0&0&\cdots\\ -\frac{2}{2}&-2&\frac{3}{2}&0&0&\cdots\\ 0&\frac{1}{2}&-2&\frac{4}{3}&0&\cdots\\ 0&0&\frac{2}{3}&-2&\frac{5}{4}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}.$ (20)

### v2

@@ 3.2 Numerical results
In this section, we gather the results of numerical simulations performed on a personal computer. The reference energies denoted $\infty$ as $M\rightarrow\infty$} correspond to the limit of ordinary space and are taken from aquino2007highly . The results provided are eigenvalues of the matrix defined in (11), shown for different numbers of layers. The energies are in hartrees and the distances in bohrs.
The large matrix values shown in Table 1 were obtained in 10 hours on a personal computer, and the code was written in C++. It is not clear whether deviations are due to numerical errors in our computation or the approximations in the reference used. We can observe that a reasonably small matrix can be used to reproduce the value of the lowest energy level accurately.
$M$ $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ reference
Table 1: Values of the ground state energy for $r_{0}=10$ and different number of layers. The largest value took approximately 1010 hours of computational time on a personal computer. The reference is taken from aquino2007highly .
@@
In tables 3 and 3, the mean value $\left\langle r\right\rangle$ is shown for various orbitals and two different cavity sizes; they are different in the size of the used matrix. A smaller matrix with the same cavity radius corresponds to a coarser underlying structure; that is, the relative size of quantum cells is larger, and the effects of quantum space are more prominent. We can observe that for small matrix sizes, the mean value of the radial coordinate of the particles increases when they are pushed outward from the centre.
We are also interested in the regime of strong effects of quantum space. We can again consider three situations with regard to the size of confinement, but now, instead of taking $\lambda\rightarrow 0$, we take $\lambda\rightarrow a_{0}$. Why is this situation interesting? Firstly, it can give some clues about the behaviour of particles in the earliest stages of the universe, where everything was dense and compressed and also energetic enough to make the effects of quantum space prominent. Secondly, it can provide some insight into gravitational collapse. Of course, in our situation, there is no gravity, but earlier studies Nicolini:2005vd suggest that the effects of quantum space prevent the formation of a perfect singularity and provide a form of outward pressure that stabilizes the core, so we can now see whether the quantumness of space would act against the squeezing force of the confinement.
While the eigenvalues of (11) give us the energy spectrum, the corresponding eigenvectors are the radial part of the solution $\Psi$. In Figure 1, we show one particular example. We can see that our solution satisfactorily matches the solution obtained in aquino2007highly , and the deviation from the unconfined solution is also prominent. We have made a similar comparison for other orbitals and values of parameters, and the matching holds for as long as we are far away from the strong quantum space regime, that is, as long as $\lambda$ remains sufficiently small.
Figure 1: Comparison of the radial part of our solution using the matrix defined (11) and the ordinary-space solution for confined hydrogen atom from reference aquino2007highly for $r_{0}=10$ and 2S orbital. We have also included the unconfined solution to show the effect of the barrier.
An important aspect of our construction is that it can be easily modified to include various terms in the potential — we have focused on the hydrogen model as on an illustrative example. To demonstrate this, we have added a linear term
$H_{\scriptsize{add}}=-5r,$ (14)
to the Hamiltonian (11) and found its spectrum and set of solutions, the first four of them are shown in Figure 2. This potential has been analysed in the context of the study of bound states of quarks using different methods in Bukor:2022xjc . This shows that the constructions have the potential to solve precisely many complicated quantum-mechanical models, at least those possessing rotational symmetry.
3.3 Mirroring states
We have observed a class of solutions that have no counterpart when the space is continuous, that is $\lambda=0$}. These have been conjectured in a previous study of the hydrogen atom in a noncommutative space pg13 ; Gal:2013 . It has been observed that in addition to states with energies $E_{n}^{I}$ that reproduce the spectrum of an unconfined hydrogen atom in ordinary space, there is an additional set of bound states for a repulsive potential of the same form. The value of these energies is
$E_{n}^{II}=\frac{2}{\lambda^{2}}-E_{n}^{I}.$ (15)
In Table 4, we compare the two energy scales obtained as smallest and largest eigenvalues of (11) with the values predicted by (15) — the match is nearly perfect. In Figure 3, we plot the radial part of the wavefunctions for an attractive potential. We obtained essentially identical results for the repulsive potential, taking (15) into account.
@@ 3.3 Mirroring states
Table 4: $E^{II}$ are the largest eigenvalues for repulsive potential, $E^{I}$ are the smallest eigenvalues for the attractive potential. The next column shows the theoretical prediction, and the last one compares this with the obtained numerical values. Energies are in Hartree.
Figure 3: Radial parts of the probability distribution for all energy eigenstates for the model (12) with $r_{0}=10$ and $M=100$. The same plot is obtained — in reverse order, starting from the largest eigenvalue, see (15) — for the same model but with the repulsive potential. Green lines show numerical solutions of (11).
A similar mirror has been observed for other choices of the potential $V(r)$. Here, we sketch the proof and underlying assumptions for the existence of mirroring spectra. That is, for a set of solutions of the eigenproblem with $V(r)\rightarrow-V(r)$ and energy spectrum (15), that satisfy:
$\left(\textbf{H}_{0}+\textbf{V}(r)\right){\cal C}$ $=E_{I}\ {\cal C},$ (16)
$\left(\textbf{H}_{0}-\textbf{V}(r)\right){\cal C}^{\prime}$ $=E_{II}\ {\cal C}^{\prime}.$
Since $\cal C$} and $\cal C$$^{\prime}$ are vectors, we can take ${\cal C}^{\prime}=\textbf{P}{\cal C}$, where P is a matrix. The next step is to multiply the first equation by P from the left and add those two equations together, obtaining
$\left(\{\textbf{P},\textbf{H}_{0}\}+[\textbf{V}(r),\textbf{P}]\right){\cal C}=\left(E_{I}+E_{II}\right)\textbf{P}{\cal C}.$ (17)
From this, we deduce two equations for P:
$\{\textbf{P},\textbf{H}_{0}\}$ $=$ $\left(E_{I}+E_{II}\right)\textbf{P}$ (18)
$[\textbf{P},\textbf{V}(r)]$ $=$ $0,$
which are solved by a diagonal matrix of the form
$\textbf{P}=\mbox{diag }\left(1,-1,1,-1,\dots\right),$ (19)
with $\left(E_{I}+E_{II}\right)=2\lambda^{-2}$, yielding (15). This matrix has a physical interpretation. The state with maximal oscillation in the fuzzy onion space has the form ${\cal C}\sim\left(1,-1,1,-1,\dots\right)$. This matrix maps the constant vector onto this, and since $\textbf{P}^{2}=\textbf{1}$, it maps the maximal oscillation vector into a constant one, a highly-oscillatory state is mapped into a nearly constant and so on; connecting the UV part of on theory into IR part of the one with flipped potential. Of some interest might be the state satisfying $\textbf{P}{\cal C}={\cal C}$. In KovacIk:2013vbk , it has been observed that while the energy spectrum is bounded by $2\lambda^{-2}$, the Heisenberg uncertainty relation vanishes for a (free) state with the energy of $\lambda^{-2}$; it might be worthwhile to investigate where this self-adjoined state has some special properties in the case with nontrivial potential as well.
3.4 Number of negative eigenvalues
In this paper, we study a finite model that has, as we have argued, a limit in which it approaches ordinary continuum three-dimensional space. Let us take the hydrogen model as an example. In the continuum limit, it has a discrete part of the energy spectrum (bounded states with negative energy) and a continuous part (scattering states with positive energy). For our finite model, the particle is always confined in the cavity, but we can still analyse how, for a fixed set of parameters, the finite number of eigenvalues split between positive and negative values, and how they behave as the continuum limit is approached. To be precise, let us fix $\lambda=0.1$ and see the energy spectrum for different values of $M$ (considering only $l=0$ states). In Figure 4, we can see that the smallest eigenvalue initially was positive and as $M$ was increased, it decreased, became negative and converged to a fixed value. Similar behaviour appeared for the rest of the eigenvalues. To summarise, by increasing the matrix size $M$, the number of eigenvalues grows and their corresponding values gradually decrease and converge to a certain value. We have computed the number of negative eigenvalues for various values of $M$ and the results are shown in Figure 5. As we can observe, their number grows proportional to $\sqrt{M}$, meaning that in the continuum limit, the ratio of the number of scattering states and bounded states grows indefinitely. There is a physical explanation for the square root growth. The average distance from the origin for a bounded state at energy level $n$ grows as $\langle r\rangle\sim n^{2}$; at least in the continuum limit with infinite volume, our model is an approximation of this. The radius of the cavity is $R=\lambda M$, and unless the state fits in it, that is $R\approx\langle r\rangle$, its wavefunction is distorted and energy increased, making it positive. Therefore, the largest negative level $n$, that is the number of bounded states, grows as $n\sim\sqrt{M}$.
Figure 4: Spectrum of the hydrogen atom with $\lambda=0.1$. For better visualisation of the eigenvalue flow, the smallest, second smallest, and so on, eigenvalues have been connected across various values of $M$.
Figure 5: Number of negative energies of the hydrogen atom with respect to M for $\lambda=1$. The blue line is fit with the function $a\sqrt{M}+b$
@@ 4 Conclusions and outlook
We have demonstrated that a simple model of quantum space can be used, in a certain limit, as a description of ordinary physics, e.g. the hydrogen atom problem. The defining feature of this model is that it is expressed in terms of matrices and can, therefore, easily be analysed on a computer.
To showcase the behaviour of the model, we have analysed the behaviour of the hydrogen atom confined in a spherical cavity and compared our results to those previously obtained in high-precision studies; the agreement is up to numerical uncertainty. We have set up a Hamiltonian expressed in terms of a matrix; its eigenvalues form the energy spectrum, and the corresponding eigenvectors for the radial dependency. The model can be easily amended to have a different form of potential, making this a rather flexible model.
In the case of the hydrogen atom, it was possible to take only a single value $(l,m)$; however, angular-dependent problems can be analyzed in the defined setting as well; one just needs to use the full form of the matrix $\Psi$ and the Hamiltonian.
In our case, we sent the constant $\lambda$ describing the quantumness of space to 0 to recover the physics in ordinary space. However, one can be motivated to choose a nonzero value to describe a space with a granular structure or to capture the essence of a finite-length scale feature of the studied model.
We have verified the behaviour of the confined particle with respect to the different radii of the cavity and showed how the particle is affected in certain regimes. We have also studied the behaviour in the strong quantum-space regime. The first important observation was that increasing the value of $\lambda$ pushes the particle away from the centre; this means that the quantumness of space manifests as an outward pressure. This is important, for example, to understand the behaviour of regularized singularity in studies of microscopic black holes sensitive to this length scale Nicolini:2005vd ; Kovacik:2021qms .
In addition to that, it has been shown that a new form of bounded states should appear for ultra-high energies pgk15 . We have shown that these states indeed appear and confirm the pattern of mirroring energies. Since the parameter, $\lambda$ can also be understood as describing a length scale that is not of quantum-space origin, but perhaps due to two granular features of the material, it will be interesting to consider whether such mirroring states might be experimentally observable. Also, our plan is to investigate the behaviour of a free particle moving in confinemed region, studying the pressure it produces (in the strong quantum space limit); those studies should provide some insight into high-density regimes of the theory of quantum space, which are needed for the description of the earliest stages or the universe or the behaviour of black hole singularities as described in pgk15 ; pg13 .
@@ Acknowledgements.
This research was supported by VEGA 1/0025/23 grant Matrix models and quantum gravity and MUNI Award for Science and Humanities funded by the Grant Agency of Masaryk University. The authors also thank the hospitality of the Physikzentrum Bad Honnef, where the work in this paper was initiated. The authors would like to acknowledge the contribution of the COST Action CA23130. We would like to thank Lukáš Konečný for his comments.
Appendix A Explicit construction of the $su(2)$ representation with Hermitian matrices of rank NN
Let us begin with the construction of the angular momentum generators satisfying
$L_{i},L_{j}]=i\varepsilon_{ijk}L_{k}.$ (20)
We can choose a basis for which eigenstates of $L_{3}$ are unit vectors of the form
$e_{l-m}=(0,\ldots,0,1,0,\ldots,0)$ (21)
such that $\hat{L}_{3}e_{m}=me_{m}$. Then, we define two matrices that increase this value by $1$, that is $L_{\pm}e_{m}\sim e_{m\pm 1}$. Taking into account the normalisation, one obtains
@@ Appendix A Explicit construction of the $su(2)$ representation with Hermitian matrices of rank NN
and $L_{-}=L_{+}^{T}$. The off-diagonal coefficients are $\sqrt{\left(l+1-m\right)m}$. Now one can take
@@ Appendix B Radial term on the fuzzy onion
${\cal{K}}=\lambda^{-2}\begin{pmatrix}-2&\frac{2}{1}&0&0&0&\cdots\\ \frac{1}{2}&-2&\frac{3}{2}&0&0&\cdots\\ 0&\frac{2}{3}&-2&\frac{4}{3}&0&\cdots\\ 0&0&\frac{3}{4}&-2&\frac{5}{4}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}$. (26)
Appendix C Numerical results for various orbitals
The following tables 5, 6, 7, 8, 9 summarize the numerical results for various orbitals with the comparison with the results obtained in aquino2007highly . After the tables, two other plots follow. Figure 6 shows a similar situation to Figure 1, while Figure 7 shows a comparison of two mirroring states.
Figure 6: Comparison of the radial part of our solution using the matrix defined (11) and the ordinary-space solution for confined hydrogen atom from reference aquino2007highly for $r_{0}=20$ and 44S orbital. We have also included the unconfined solution to show the effect of the barrier.
Figure 7: Comparison of the eigenvectors for the third smallest eigenvalue in the case of attractive potential and the third largest eigenvalue for the repulsive potential; these states have dual energy according to (15). These values were obtained numerically for $M=1000$, $r_{0}=40$ and $n=3$.
[…]
Table 5: The 1S orbital.
[…]
20 -0.124987114312918 -0.124677720985899 -0.124984183578311 -0.124987102906836
Table 6: The 2S orbital.
[…]
Table 7: The 3S orbital.
[…]
Table 8: The 2P orbital.
[…]
Table 9: The 3P orbital.